IONIC REACTIVITY OF SMALL GAS PHASE MOLECULAR CLUSTERS

\underline{NICOLA M TONGE}, EWAN C Mac MAHON and MARTIN C R COCKETT, University of York, Department of Chemistry, Heslington, YO10 5DD, United Kingdom.

The study of ionised molecular clusters provides an insight into ion-molecule chemistry; ubiquitous in atmospheric, flame and plasma processes. The neutral fluorobenzene-ammonia system, essentially unreactive in the gas phase, will react in solution under high temperature and pressure conditions. However, fluorobenzene^{+} will react, via nucleophilic substitution to form aniline^{+} and HF. In combination with the aquisition of REMPI spectra, we have observed the following reactions, post-ionisation, of fluorobenzene^{+}-(NH_{3})_{n} (n=1,2) clusters:

C_{6}H_{5}F^{+}.(NH_{3})_{2} ⟶ C_{6}H_{5}NH_{2}^{+}

C_{6}H_{5}F^{+}.NH_{3} ⟶ NH_{3}^{+} + C_{6}H_{5}F

While nucleophilic substitution satisfactorily accounts for the formation of aniline^{+}, it is more difficult to rationalise the formation of NH_{3}^{+} on thermodynamic grounds. By performing ab initio calculations we hope to be able to postulate a mechanism for this reaction.